OpenStudy (monkey*):

Find the minimum value of f(x,y)=y-x+1 for the polygonal convex set determined by this system of inequalities. X> or = 0 Y> or =0 2x+Y< or =4

OpenStudy (monkey*):

Any help?

OpenStudy (irishboy123):

\(f(0,0) = 1\) \( f(x,y)=y-x+1\) \(f_x = -1, \, f_y = 1\) so \(f_{min} = f(2,0) = -1\)

OpenStudy (monkey*):

Would that be my answer? @IrishBoy123

OpenStudy (irishboy123):

that would be my answer, and if it is wrong, i could justify my approach it is not clear to me how "you" are supposed to approach this, eg what are you learning right now. i just blagged it using some general knowledge if you want more help i can tag some poeple

OpenStudy (monkey*):

Yes, please. I'm completely lost on this problem.

OpenStudy (irishboy123):

before i do so, what are you learning now? that might provide a clue as to how you are expected to approach/solve it

OpenStudy (monkey*):

Would your answer be (2,0) or -1? Because if it's the (2,0), that would make more sense to me. But I'm learning system of linear equations and inequalities. Matrices also.

OpenStudy (irishboy123):

it is 2,0 i have looked on line and you are supposed to plot the region and then just plug in values for x and y. that kinda is what the calculus shortens so 

OpenStudy (irishboy123):